# egoPPG: Heart Rate Estimation from Eye-Tracking Cameras in Egocentric Systems to Benefit Downstream Vision Tasks

### v1

@@ 5.3 Experiments setup
5.3.1 Training
We trained all implemented models using five-fold cross-validation split by participants to ensure a strict separation between training, validation, and test sets. We iteratively held out the data from five participants (20%) as the test set, two as validation, and used the remaining participants as the training set. The training was conducted with a batch size of 4 for 100 epochs, a learning rate of 0.0009, and mean squared error (MSE) as the loss function. In addition to our model, we used four baseline networks (one state-space and three CNNs), which have state-of-the-art performance for rPPG, to compare the performance of our proposed model to the performance of these established models. Our model was trained on a GeForce RTX 4090, with a total runtime of about 15 hours for all folds.
5.3.2 Metrics
To assess model accuracy, we use the mean absolute error (MAE), root mean squared error (RMSE), mean absolute percentage error (MAPE), and Pearson correlation (r) using a non-overlapping 60-second sliding window, as commonly used for rPPG [18, 92, 48, 55, 49].
5.3.3 Video sampling rate
While we recorded the eye tracking videos with 30 fps, large-scale datasets such as EgoExo4D [33] or Nymeria [51] used only 10 fps. To assess the impact of reduced frame rates, we evaluated i) model performance when downsampling our videos to 10 fps by retaining only every third frame and ii) model performance when downsampling to 10 fps, then linearly interpolating between frames to upsample to 30 fps again. For both scenarios, we train from random initialization.
5.3.4 Hyperparameter optimization
We also provide insights into optimal hyperparameter configurations by analyzing the impact of image size, input length, loss function, and choice of label (see Tab. 9).
@@ 6 Downstream use for proficiency estimation
To demonstrate the utility of predicting a user’s physiological state for egocentric vision applications, we use the user proficiency estimation benchmark from the EgoExo4D dataset, which contains over 5000 videos from 740 participants performing skilled human activities [33]. The user proficiency estimation benchmark aims to classify the proficiency of a user (novice, early expert, intermediate expert, late expert) using only egocentric video clips (Ego), only exocentric video clips (Exo), or all video clips together (Ego + Exo). Our goal was to assess if we can improve the performance of the current baseline model (TimeSFormer [6]) when integrating our predicted HR data into the network. This results in two additional configurations: using egocentric videos and HR together (Ego + HR), and using all video clips and HR together (Ego + Exo + HR). To predict the continuous HRs for all EgoExo4D videos, we use our proposed method, pre-trained on egoPPG-DB.
We implement the TimeSFormer model in exactly the same configuration as for the current benchmark results [33] with a clip size of 16 frames and a sampling rate of 16, trained for 15 epochs on four GeForce RTX 4090. We use all videos of the EgoExo4D dataset, for which the proficiency estimation labels are available (using the official benchmark training and validation sets) and which have at least 16 frames at a sampling rate of 16, resulting in totally 2044 videos. From the official training set, we use 10% as validation, and the held-out official validation set for testing. We summarize our predicted HR data by calculating five features (mean, standard deviation, minimum and maximum HR, and mean HR change) for the corresponding videos. To integrate our HR predictions into the TimeSFormer architecture, we simply feed our calculated features into a 50-parameter linear layer and concatenate the output with the backbone’s output before feeding it into the network head. We train all models from random initialization and evaluate proficiency estimation using top-1 accuracy per the EgoExo4D protocol.
@@ 7.1 Heart rate estimation
7.1.1 Baseline
We employed signal processing to verify that the BVP signal is present in the eye tracking videos, to determine in which regions the SNR is highest, and to establish a baseline for comparison. Since the glasses remain mostly stable throughout the recording, we manually define two spatial cropping regions per participant. One region that includes mostly skin, and one region that includes mainly eyes (see Fig. 4). After spatially cropping the images, we calculate the mean pixel intensity values and remove motion artifacts by discarding any changes in the signal that are outside the first and third quantiles of the signal’s changes. Finally, we filter the signal with a 4th order Butterworth bandpass filter between 0.6 and 3.0 Hz (corresponding to 36 to 180 bpm) to obtain the BVP (see Fig. 6).
7.1.2 EgoPulseFormer: an egoPPG method
Using our proposed network, we obtain an MAE of 8.82 bpm and a correlation of 0.81 between our predicted HR and the ground truth HR (see Tab. 2). This is an improvement of 3.27 bpm (27.0%) of the MAE and 0.15 for the correlation compared to the second-best network (PhysNet [92]). Split by activity, we obtain the lowest MAE while the participants are watching a video (MAE of 6.25 bpm) and the highest MAE during exercising on a bike (MAE of 13.51 bpm) and while dancing (MAE of 10.02 bpm), which are both also the tasks with the highest normalized motion magnitude. We define the motion magnitude as the root mean squared sum of the absolute differences across the 3-axis IMU recorded by the Aria glasses. We, then, normalize it between zero and one across all activities to get a measure of the amount of motion of each activity. Furthermore, using simple signal processing and spatial cropping, we obtain an MAE of 12.40 and a correlation of 0.50 when using the skin region around the eyes. When using the eyes regions as input, the MAE increases to 14.60, and the correlation drops to 0.20. This is also reflected in spatial attention maps that our model implicitly learns (see Fig. 4), which exclude the eyes for predicting the HR. To qualitatively cross-check these results, Fig. 6 shows an example plot of the raw mean intensity values (before filtering) of the skin region compared to the eye region with the BVP clearly visible for the skin region. Tab. 4 shows the results when downsampling our videos to 10 fps. The MAE increases to 12.40 bpm and the correlation decreases to 0.52 when training and testing using 10 fps. When upsampling the videos again to 30 fps using linear interpolation, the MAE decreases to 10.20 bpm and the correlation increases to 0.77.
@@ 7.1.2 EgoPulseFormer: an egoPPG method
Mean intensity eyes 14.60 18.18 18.37 0.20
@@
Mean intensity skin 12.40 15.54 15.29 0.50
@@
EgoPulseFormer (ours) 8.82 12.03 10.82 0.81
Improvement over -3.27 -3.4 -4.32 +0.15
@@
Table 2: Results for HR prediction from eye tracking videos using different models (EgoPulseFormer and established rPPG baselines).
Activity $\mu$ HR Motion MAE RMSE MAPE
magnitude
Video 71.45 0 6.25 8.50 9.88
Office 75.65 0.45 8.90 11.97 12.60
Kitchen 85.31 0.54 8.75 11.11 10.42
Dancing 89.07 1.00 10.02 12.79 11.22
Bike 113.06 0.77 13.51 16.52 11.04
Walking 93.71 0.30 7.02 9.31 6.70
Table 3: Results for HR prediction split by activity.
@@
10 fps (other datasets) 12.40 16.80 14.10 0.52
Upsampled to 30 fps 10.20 13.56 12.78 0.77
@@ 7.2 Downstream task: proficiency estimation
Tab. 5 summarizes the results of our experiments to evaluate the value of HR estimation for the proficiency estimation benchmark on the EgoExo4D dataset. We see that integrating our predicted HRs into the TimeSFormer model [6] improved accuracy for all scenarios but one. Additionally, we also achieved the highest accuracy for each individual scenario with our HR integration for the same scenarios. When combining the egocentric videos with our predicted HRs, we achieved an overall accuracy of 45.20%, a 14.1% increase compared to using egocentric videos alone. The largest gains appeared in the cooking and dancing tasks, where accuracy rose from 20.00% to 40.00% and from 43.44% to 53.27%, respectively. Also, when using the egocentric videos, exocentric videos, and our predicted HRs together, the accuracy increased by 8.64% from 39.00% to 42.37% compared to using only the egocentric and exocentric videos. Using only exocentric videos yielded the lowest overall accuracy at 35.93%.
@@
Basketball 38.00 45.45 47.47 49.24 49.24 49.49 52.52
Cooking 0.00 20.00 40.00 33.75 40.00 25.00 40.00
Dancing 24.59 43.44 53.27 45.08 48.36 50.82 55.73
@@
Bouldering 15.29 24.50 27.81 10.26 12.58 15.89 18.54
Soccer 62.50 50.00 56.25 76.56 75.00 75.0 62.50
Overall 27.80 39.69 45.29 35.93 35.93 39.00 42.37
Table 5: Results for proficiency estimation benchmark on EgoExo4D dataset. Note that for all scenarios except Soccer, the accuracy increases when integrating EgoPulseFormer’s heart rate estimate into the existing and otherwise unmodified baseline model.
@@ 8.1 Heart rate estimation
When evaluating our proposed method on egoPPG-DB, we showed that HR can reliably be predicted from eye tracking videos of unmodified egocentric vision headsets. Compared to the performances achieved on popular rPPG datasets such as PURE [78], UBFC-RPPG [7], and MMPD [80], we achieved very competitive performance in challenging settings [49]. While the lowest reported MAE for UBFC-RPPG in the rPPG-toolbox [49] is 1.21 bpm, its participants sit almost motion-free with their eyes closed. Already on MMPD, a dataset with varied lighting and little motion, the lowest reported MAE increases to 10.23 bpm [49], as it incorporates more realistic challenges due to recordings on mobile devices with light motion (head rotation, talking, and taking selfies). In comparison, using EgoPulseFormer, we achieved a lower MAE of 8.82 bpm while even having tasks with heavy motion (dancing, walking stairs) and very high HR changes (44–164 bpm).
@@ 8.1.1 Performance depending on activity
Analyzing our results split by activity (see Tab. 3), we see that the tasks with the highest mean HR (bike, 113 bpm) and motion magnitude (dancing) also have the highest MAE. Given the substantially bigger motion artifacts and HR variability (HRs between 44 and 164 bpm) in our dataset — captured during diverse everyday activities and physical exercises — we believe that our results demonstrate the robustness of EgoPulseFormer in more dynamic, everyday conditions.
8.1.2 Performance depending on method
For our model, we leveraged spatial attention maps to improve performance. When qualitatively analyzing the learned spatial attention maps, we see that our model implicitly learned to exclude the eyes for estimating the BVP from the eye tracking videos (see Fig. 4). This aligns with our obtained qualitative (see Fig. 6) and quantitative (see Tab. 2) results using simple signal processing, which show a higher SNR for the skin region compared to the eyes. When analyzing the performance of other state-of-the-art rPPG models on our dataset, we found that only PhysNet [92] and PhysMamba [90] achieved reasonable performances, though their results did not reach our performance.
@@ 8.2 Benefits for proficiency estimation downstream
We found that incorporating HR data into the baseline model of the proficiency estimation task substantially improved accuracy across both configurations. The egocentric videos combined with the HR achieved the highest overall accuracy at 45.29%, marking a 14.1% increase over using only egocentric videos (39.69%). Adding HR especially improved accuracy for cooking (from 20% to 40%) and dancing (from 43.4% to 53.3%), which had the lowest accuracies besides bouldering when using only egocentric videos, demonstrating the value of HR in enhancing model performance. Combining egocentric videos, exocentric videos, and HR provided further accuracy gains for some scenarios, achieving the best results for basketball, cooking, and dancing. Results using exocentric views alone were lower overall, which is consistent with benchmark results [33]. These findings suggest that adding HR data as an auxiliary signal enhances performance for the proficiency estimation benchmark on the EgoExo4D dataset. For training and testing, we used the available subset of EgoExo4D videos for which proficiency labels are available, following the official training and validation splits. While our used data shows slight variations from the official release in majority class distributions and accuracy scores, the observed trends align well with the established benchmark results.
8.3 Limitations and future work
We observed the highest MAE in tasks with elevated HRs and motion, such as dancing and biking. We believe that promising approaches to address these limitations could be to record more tasks with high HRs and integrating the IMU data from the glasses. Furthermore, future studies could extend data collection to outdoor settings to assess the impact of varying lighting conditions. While we ensured a balanced gender ratio (13 male, 12 female), our sample size of 25 participants restricts broader demographic conclusions. In future work, we aim to expand the dataset to investigate performance variations across different age groups, skin types, and ethnicities. Finally, we believe that it is a highly interesting problem to explore further downstream applications of a user’s physiological state, such as personalized feedback, autonomous assistance, as well as health-related applications.
@@ 9 Conclusion
We introduced egoPPG, a novel task for egocentric vision systems to extract the wearer’s heart rate for integrating their physiological state into egocentric vision tasks downstream. Our method EgoPulseFormer processes input from the eye tracking cameras on unmodified egocentric vision systems to robustly estimate the person’s HR in various everyday scenarios. We validate EgoPulseFormer’s robustness on our dataset egoPPG-DB and demonstrate significant improvements over existing rPPG models. With HR estimations from EgoPulseFormer we significantly improve the proficiency estimation benchmark on the large-scale EgoExo4D dataset. Our results emphasize the potential of physiological insights obtained via egoPPG methods for further egocentric vision applications.

### v2

@@ 5.3 Experiments setup
Training. We trained all models using five-fold cross-validation split by participants to ensure a strict separation between training, validation, and test sets. We iteratively held out the data from five participants (20%) as the test set, two as validation, and trained on the remaining with a batch size of 4 for 100 epochs, a learning rate of 0.0009, and mean squared error (MSE) as loss. In addition to our model, we used ten state-of-the-art rPPG baseline networks to compare the performance of our proposed model to these established models. Our model was trained on a GeForce RTX 4090, with a total runtime of about 20 hours for all folds.
Evaluation. To calculate the HR, we filter the predicted BVP with a Butterworth filter (0.5–2.8 Hz) and then detect peaks. To assess model accuracy, we use the mean absolute error (MAE), root mean squared error (RMSE), mean absolute percentage error (MAPE), and Pearson correlation (r) using a non-overlapping 60-second window [16, 45, 46].
Video sampling rate. While we recorded the eye tracking videos with 30 fps, large-scale datasets such as EgoExo4D [27] or Nymeria [50] used only 10 fps. To assess the impact of reduced fps, we evaluated model performance when (1) downsampling our videos to 10 fps by retaining only every third frame and (2) downsampling to 10 fps, then linearly interpolating between frames to upsample to 30 fps.
@@ 6 Downstream use for proficiency estimation
To demonstrate the utility of predicting a user’s physiological state for egocentric vision applications, we use the user proficiency estimation benchmark from the EgoExo4D dataset, which contains over 5000 videos from 740 participants performing skilled human activities [27]. This benchmark aims to classify the proficiency of a user (novice, early expert, intermediate expert, late expert) using only egocentric videos (Ego), only exocentric videos (Exo), or all videos together (Ego + Exo). Our goal was to assess if we can improve the performance of the current baseline model (TimeSFormer [5]) when integrating our predicted HR data into the network. This results in three additional configurations: using egocentric/exocentric videos and HR (Ego + HR/Exo + HR) and using all videos and HR (Ego + Exo + HR). To predict the continuous HRs for all EgoExo4D videos, we use PulseFormer, pre-trained on egoPPG-DB.
We implement the TimeSFormer model in exactly the same configuration as for the benchmark results [27] with a clip size of 16 frames and a sampling rate of 16, trained for 15 epochs. We use all videos of the EgoExo4D dataset, for which the proficiency estimation labels are available (using the official benchmark training/validation sets) and which have at least 16 frames at a sampling rate of 16, resulting in 2044 videos. From the official training set, we use 10% as validation, and the held-out official validation set for testing. We summarize our predicted HR data by calculating five features (mean, STD, minimum and maximum HR, and mean HR change) for the corresponding videos. We integrate these features via normalization and a 50-parameter linear layer whose output we concatenate with the output of TimeSFormer’s backbone before feeding it into the classification head. We train all models from random initialization and evaluate using top-1 accuracy per EgoExo4D protocol.
@@ 7.1 Heart rate estimation
7.1.1 Signal-processing baseline
We employed signal processing to verify that the BVP signal is present in the eye tracking videos, to determine in which regions the SNR is highest, and to establish a baseline (see Tab. 2). Since the glasses remain mostly stable throughout the recording, we manually define two spatial cropping regions per participant. One region that includes mostly skin, and one region that includes mainly eyes (see Fig. 4). We calculate the mean pixel intensity for both regions, remove motion artifacts by discarding any changes outside the interquartile range and finally filter the signal with a 4th order Butterworth bandpass filter (0.5 to 2.8 Hz) to obtain the BVP (see Fig. 5 in Supplementary).
7.1.2 PulseFormer method
Using our proposed network PulseFormer, we obtain an MAE of 7.67 bpm and a correlation of 0.85 between our predicted HR and the ground truth HR (see Tab. 2). This is an improvement of 2.40 bpm (23.8%) of the MAE and 0.13 for the correlation compared to the current SOTA FactorizePhys [38]. Split by activity, we obtain the lowest MAE while the participants are watching a video (MAE=5.52 bpm) and the highest MAE during exercising on a bike (MAE=12.91 bpm), which is the task with the highest mean HR (113.1 bpm) and the second highest motion magnitude. In addition, the MAE decreases for all activities when adding MITA, with the greatest performance improvement for dancing. We define the motion magnitude as the root-mean-squared sum of the absolute differences across the 3-axis IMU recorded by the Aria glasses and normalize it between zero and one across all activities to get a measure of motion of each activity. See Fig. 6 (Supplementary) for a boxplot of the MAEs of PulseFormer’s predictions. Using signal processing, we obtain an MAE of 12.40 bpm when using the skin region around the eyes and an MAE of 14.60 bpm using the eye regions as input. This is also reflected in the spatial attention maps that our model implicitly learns (see Fig. 4), which exclude the eyes to predict the HR. To qualitatively cross-check these results, Fig. 5 (Supplementary) shows an example plot of the raw mean intensity values (before filtering) of the skin region compared to the eye region, with the BVP clearly visible for the skin region. Tab. 4 shows the results when downsampling our videos to 10 fps. MAE increases to 11.13 bpm and the correlation decreases to 0.7 when training and testing using 10 fps. When upsampling the videos again to 30 fps, the MAE decreases to 10.20 bpm and the correlation increases to 0.77. In Sec. 16 (Supplementary), we show that PulseFormer also outperforms the baselines in a cross-dataset evaluation.
@@ 7.1.2 PulseFormer method
ContrastPhys+ [81] 19.12 24.13 22.57 0.21
RhythmMamba [99] 15.05 19.78 17.46 -0.16
Baseline eyes 14.60 18.18 18.37 0.20
@@
RhythmFormer [98] 13.13 17.43 14.73 0.51
Baseline skin 12.40 15.54 15.29 0.50
@@
PhysFormer [93] 10.71 13.97 12.69 0.72
PulseFormer w/o SA 10.49 13.62 12.83 0.73
FactorizePhys [38] 10.07 13.43 12.36 0.67
PulseFormer w/o MITA 8.82 12.03 10.82 0.81
PulseFormer (ours) 7.67 10.69 9.45 0.85
Improvement over -2.40 -2.74 -2.91 +0.13
@@
Table 2: Results for HR prediction from eye tracking videos using different models (PulseFormer, PulseFormer without SA, PulseFormer without MITA and established rPPG baselines).
Activity $\mu$ HR Motion PulseFormer PulseFormer
magnitude w/o MITA
Video 71.5 0 5.52 5.97
Office 75.7 0.45 7.50 8.22
Kitchen 85.3 0.54 7.22 8.89
Dancing 89.1 1.00 7.85 10.54
Bike 113.1 0.77 12.91 14.62
Walking 93.7 0.30 8.23 8.29
Table 3: Results for HR prediction (MAE) split by activity using PulseFormer and PulseFormer without MITA.
@@
10 fps (other datasets) 11.13 15.18 12.28 0.70
Upsampled to 30 fps 10.18 13.07 12.48 0.77
@@ 7.2 Downstream task: proficiency estimation
Tab. 5 summarizes the results of our experiments to evaluate the value of HR estimation for the proficiency estimation benchmark on EgoExo4D. We see that integrating our predicted HRs into the TimeSFormer model [5] improved accuracy for all scenarios but one (Soccer). We also achieved the highest accuracy for each of these individual scenarios with our HR integration. When combining the egocentric videos with our predicted HRs, we achieved an overall accuracy of 45.29%, a 14.1% increase compared to using egocentric videos alone. The largest gains appeared in the cooking and dancing tasks, where accuracy rose from 20.00% to 40.00% and from 43.44% to 53.27%. Also, when using the egocentric and exocentric videos, and our predicted HRs together, the accuracy increased by 12.67% (4.94 percentage points) from 39.00% to 43.94% compared to using only the egocentric and exocentric videos.
@@
Basketball 38.00 45.45 47.47 48.48 48.48 49.49 50.50
Cooking 0.00 20.00 40.00 35.00 40.00 25.00 40.00
Dancing 24.59 43.44 53.27 42.62 48.36 50.82 59.84
@@
Bouldering 15.29 24.50 27.81 8.61 12.58 15.89 21.19
Soccer 62.50 50.00 56.25 81.25 75.00 75.00 62.50
Overall 27.80 39.69 45.29 34.75 37.67 39.00 43.94
Table 5: Results for proficiency estimation benchmark on EgoExo4D dataset. Note that for all scenarios except Soccer, the accuracy increases when integrating PulseFormer’s heart rate estimate into the existing and otherwise unmodified baseline model.
@@ 8.1 Heart rate estimation
Evaluating PulseFormer on egoPPG-DB, we showed that HR can be reliably predicted from eye tracking videos and IMU signals from unmodified egocentric vision headsets. While SOTA rPPG models (e.g., PhysFormer) achieve MAEs as low as 0.50 bpm [93, 98] on datasets such as UBFC-RPPG [6] or OBF [44], these datasets captured participants while calmly sitting at a table looking at the camera (with very little motion or HR changes). However, during even just light motion (e.g., on MMPD [82] or VIPL-HR [63]), their MAE increases to 5.0–12.0 bpm [93, 98].
On egoPPG-DB, which contains much stronger motion (dancing, exercise bike) and HR fluctuations (between 44–164 bpm), PulseFormer’s MAE is 7.67 bpm and outperforms the rPPG SOTA FactorizePhys (MAE=10.07 bpm). Given the strong motion and diverse everyday activities in egoPPG-DB, we believe that our results demonstrate the robustness of PulseFormer in dynamic, everyday conditions. For context, even HR estimates from contact sensors tightened to the body (e.g., Apple Watch) yield an MAE of 3.0 bpm during rest and an MAE of 4.6 bpm on a bike [24].
8.1.1 Performance depending on method
We introduced MITA and leverage SA modules to improve the performance of our model. The performance of PulseFormer decreases from 7.67 bpm to 8.82 bpm when removing the MITA and to 10.49 bpm when removing the SA modules (see Tab. 2). When qualitatively analyzing the learned SA maps, we see that our model implicitly learned to exclude the eyes for estimating BVP from the eye tracking videos (see Fig. 4). This aligns with our results using signal processing, obtaining better performance for the skin region compared to the eyes (see Tab. 2).
@@ 8.1.2 Performance depending on activity
Analyzing our results split by activity (see Tab. 3), we obtain the highest MAE when exercising on a bike (MAE=12.91 bpm) and the lowest MAE when watching a video (MAE=5.52 bpm). While watching a video yields the lowest MAE, it is higher than MAEs typically reported for rPPG datasets, such as UBFC-RPPG [6], despite similar levels of motion. We attribute this to two factors: first, the higher variability in HR across egoPPG-DB, requiring improved generalization, and second, the inherent motion artifacts in eye tracking videos from blinking and natural eye movements, even during static tasks like watching videos. Such inherent motion artifacts and, e.g., slipping glasses can make capturing rPPG more difficult in this manner. Furthermore, although dancing has the highest motion magnitude, its MAE (7.85 bpm) is comparable to that of lower-motion tasks such as office and kitchen activities. When comparing performance with and without our MITA module, we observe an improvement of 2.6 bpm for dancing, indicating that MITA effectively addresses motion-induced artifacts.
@@ 8.2 Benefits for proficiency estimation downstream
We found that incorporating HR data into the baseline model of the proficiency estimation task substantially improved accuracy across all three configurations. The egocentric videos combined with the HR achieved the highest overall accuracy at 45.29%, marking a 14.1% increase over using only egocentric videos (39.69%). Adding HR especially improved accuracy for cooking (from 20% to 40%) and dancing (from 43.44% to 53.27%), which had the lowest accuracies besides bouldering when using only egocentric videos, demonstrating the value of HR in enhancing model performance. Combining egocentric videos, exocentric videos, and HR provided further accuracy gains for some scenarios, achieving the best results for basketball, cooking, and dancing. Results using exocentric views alone were lower overall, which is consistent with benchmark results [27]. Soccer was the only scenario, for which the performance decreased for Exo+HR and Ego+Exo+HR. We see two reasons for that. 1) Of EgoExo4D’s 2044 official train/test videos, only 77 are soccer, making it the scenario with the least training/test data by far. 2) Our HR estimates may be less accurate for “Stop-and-Go” sports, which are not captured in egoPPG-DB right now. In Tab. 9 (Supplementary), we show that we obtain the best downstream performance when using the HR features calculated with PulseFormer compared to the baselines. For training and testing, we used the available subset of EgoExo4D videos for which proficiency labels are available, following the official training and test splits. While our used data shows slight variations from the official release in majority class distributions and accuracy scores, the observed trends align well with the established benchmark results.
8.3 Broader impacts
Beyond health applications, such as predicting stress and fatigue [65, 60, 9], cardiac measurements could also help models better understand user behavior to, e.g., improve personalized assistance [42]. Furthermore, we believe that our approach requires the user’s explicit consent, regardless of application. Mechanisms must make users aware of measurements and require consent, e.g., on the Aria platform.
@@ 9 Conclusion
egoPPG is a novel task for egocentric vision systems to extract the wearer’s heart rate for integrating their physiological state into egocentric vision tasks downstream. We have introduced PulseFormer, a method that processes input from the eye tracking cameras on unmodified egocentric vision systems and fuses them with motion cues from the headset’s IMU to robustly estimate the person’s HR in various everyday scenarios. We validate PulseFormer’s robustness on our dataset egoPPG-DB and demonstrate significant improvements over existing rPPG models. With HR estimations from PulseFormer, we also significantly improve the proficiency estimation benchmark on the large-scale EgoExo4D dataset. Our results emphasize the potential of physiological insights obtained via egoPPG methods for further egocentric vision applications. By making our dataset available to the community, we aim to support physiological state estimation via HR in future research and new downstream tasks for egocentric vision systems. Given our promising results, we believe that future work could now focus on collecting more participants with a broader demographic background across different age groups, skin types, and ethnicities, and also extend data collection to outdoor settings to assess the impact of varying lighting conditions.